OpenStudy (anonymous):

Solve the equation a/a^2-64+3/a-8=2/a+8

pooja195 (pooja195):

−72a^2+4a=8a^2+2a(Simplify both sides of the equation) −72a2+4a−(8a^2+2a)=8a^2+2a−(8a^2+2a)(Subtract 8a^2+2a from both sides)

OpenStudy (anonymous):

i doubt it

pooja195 (pooja195):

−80a^2+2a=0 Factor that.

pooja195 (pooja195):

@satellite73 then maybe you should do it.

OpenStudy (anonymous):

\[\frac{a}{a^2-64}+\frac{3}{a-8}=\frac{2}{a+8}\] is my guess

OpenStudy (anonymous):

course @pooja195 i could be wrong

pooja195 (pooja195):

\[\huge~\rm~\frac{ a }{ a }^2-64+\frac{ 3 }{ a-8}=\frac{ 2 }{ a+8 }\] I was thinking this :/ (sorry for the attitude lol)

OpenStudy (triciaal):

@maysum can you draw to verify?

OpenStudy (triciaal):

my guess would be like @satellite73 difference of squares to be applied

OpenStudy (anonymous):

I think I am supposed to find two answers at the end, negative and positive?

OpenStudy (triciaal):

before we can get to the end can you rewrite the problem in the draw or using parenthesis?

OpenStudy (triciaal):